A large lump of zinc is placed in 100 gram of solution of HCL. 150 ml of hydrogen at 15° C and 740 mm of Hg are evolved. calculate the percentage of HCl in the solution.

Grade:

1 Answers

Shweta Ahire
5 years ago
Use PV= NRT, for H2
740/760 × 150/1000 = n × 0.0821 × 288
n = Moles of H2 evolved
So n = 0.00618
Moles of HCl is same as the is half of the mole of H2 because HCl contain only H 
So moles of HCl = 0.00309
Mass of HCl = 0.00309× 36.5 =  0.11 gm
 
So percentage of HCl = 0.11/100  × 100 = 0.11%